# Separable commutative algebras in equivariant homotopy theory

### v1

@@ 11.1. Free $G$-spectra
Consider $R\in\operatorname{CAlg}(\operatorname{Sp}_{G})$ such that $R\otimes\widetilde{E}G\simeq 0$ and $\Phi^{e}(R)$ separably closed. An example to keep in mind is Atiyah’s K-theory with reality where $G=C_{2}$, see [BCH+23, Example 8.15]; note that $\Phi^{e}KR=KU$ is separably closed by [NP24, Theorem 12.9]. Consider the pullback square from Theorem 8.2 with $R$ as above, $\mathcal{F}=\varnothing$ and $K=e$. The right top vertex of the square is
@@ 11.2. Borel $G$-spectra
If $R$ is separably closed, so that $\operatorname{CAlg}^{\operatorname{sep}}(\operatorname{Perf}(R))\simeq\operatorname{Fin}^{\operatorname{op}}$, then the above functor is also essentially surjective since $\underline{R}\otimes\mathbb{D}(\Sigma_{+}^{\infty}X)$ is a separable commutative algebra in $\operatorname{Mod}_{\operatorname{Sp}_{G}}(\underline{R})^{\omega}$ hitting the finite finite $G$-set $X$.

### v2

@@ 12.1. Free $G$-spectra
Consider $R\in\operatorname{CAlg}(\operatorname{Sp}_{G})$ such that $R\otimes\widetilde{E}G\simeq 0$ and $\Phi^{e}(R)$ separably closed. An example to keep in mind is Atiyah’s K-theory with reality where $G=C_{2}$, see [BCH+25, Example 8.15]; note that $\Phi^{e}KR=KU$ is separably closed by [NP24, Theorem 12.9]. Consider the pullback square from Theorem 8.2 with $R$ as above, $\mathcal{F}=\varnothing$ and $K=e$. The right top vertex of the square is
@@ 12.2. Borel $G$-spectra
If $R$ is separably closed, so that $\operatorname{CAlg}^{\operatorname{sep}}(\operatorname{Perf}(R))\simeq\operatorname{Fin}^{\operatorname{op}}$, then the above functor is also essentially surjective since $\underline{R}\otimes\mathbb{D}(\Sigma_{+}^{\infty}X)$ is a separable commutative algebra in $\operatorname{Mod}_{\operatorname{Sp}_{G}}(\underline{R})^{\omega}$ hitting the finite $G$-set $X$.